# Sawtooth crash in tokamak as a sequence of Multi-region Relaxed MHD equilibria

### v1

@@ I Introduction
The main purpose of the current paper is to examine the suitability of MRxMHD in modelling a sawtooth crash. Sawtooth is a well-known relaxation process in tokamaks von Goeler, Stodiek, and Sauthoff (1974) and current-carying stellarators Nagayama et al. (2003); Zanini et al. (2020). It is triggered when a $q=1$ surface presents inside the plasma volume with the temperature profile undergoing periodic crashes and flattenings, leading to worse confinement in general. Here $q$ denotes the safety factor. The crash is usually followed by a ramp-up recovering phase during which the plasma internal energy re-accumulates. The physical mechanism of the sawtooth is still under debate Chapman (2011). The two most cited models are the Kadomstev model Kadomtsev (1975) and the Wesson model Wesson (1986, 1990). The Kadomstev model attributes the crash to a $m=n=1$ internal kink mode and magnetic reconnection, where $m$ and $n$ are the poloidal and toroidal mode numbers, respectively. During a crash, flux surfaces and their neighbourhood with the same helical flux label $\chi=q_{\text{res}}\psi_{p}-\psi_{t}$ reconnects, conserving $\chi$ itself and the differential toroidal flux. Here $\psi_{p}$ and $\psi_{t}$ are the poloidal flux and toroidal flux, respectively, while $q_{\text{res}}=1$ is the resonant safety factor. The Wesson model, on the other hand, relies on interchange instabilities, which are more prominent at a relatively high plasma $\beta$, where $\beta=2\mu_{0}p/B^{2}$ is the ratio between the plasma kinetic energy and magnetic energy, with $\mu_{0}$ being the vacuum magnetic permeability, $p$ the plasma pressure, and $B$ the magnetic field strength.
@@ II The MRxMHD model
[…]
@@
×$\nabla\times\bm{B}=\mu_{i}\bm{B},\quad\text{within }\mathcal{R}_{i},$ (4)
@@
[…]
where ${\hat{\bm{n}}}$ is the normal unit vector on the interfaces. In the view of the variational principle presented here, $\mu_{i}$ is a dependent variable determined by matching the helicity constraint. Extremising (3) with respect to the location of $\mathcal{I}_{i}$ leads to the force balance condition on each interface given by
@@
[…]
@@ III.1 Equilibrium setup
A linear resistive MHD stability calculation with PHONIEX Blokland et al. (2007) shows the equilibrium is unstable to tearing/resistive-kink mode depending on the resistivity. This is demonstrated in FIG. 3 where the linear growth rate $\gamma$ is plotted as a function of plasma resistivity $\eta$. The growth rate $\gamma$ is normalised to the Alfvén frequency on axis $\omega_{A0}=V_{A0}/R_{0}=B_{0}/(R_{0}\sqrt{\mu_{0}\rho_{0}})$, where $\rho_{0}$ is the density on axis and $\mu_{0}$ the plasma vacuum permeability, while the resistivity $\eta$ is in the unit of $\mu_{0}\omega_{A0}R_{0}^{2}$, i.e. the inverse magnetic Reynolds number. At low resistivity, the growth rate follows the scaling $\gamma\sim\eta^{3/5}$, a typical relationship for a tearing mode. At a higher resistivity, it gradually transits into a resistive kink mode Coppi et al. (1976) with scaling $\gamma\sim\eta^{1/3}$. Note that the equilibrium is stable to the ideal internal kink mode Bussac et al. (1975) as no instability is found in the limit $\eta\rightarrow 0$ since $\beta=0$, and the cross-section circular.
@@ III.2 Initial MRxMHD equilibrium
Within the reconnection region, we also need to set up a few interfaces. Again, according to the Kadomtsev model, flux surfaces with the same helical flux will reconnect. We therefore place pairs of interfaces at equal helical fluxes based on the VMEC equilibrium. This is demonstrated by FIG. 4. Starting from the subregion containing the $q=1$ surface, we put $\mathcal{I}_{-1}$ inside the $q=1$ surface and $\mathcal{I}_{1}$ outside, such that $\chi(\mathcal{I}_{-1})=\chi(\mathcal{I}_{1})$. Similarly, we place $\mathcal{I}_{-2}$ and $\mathcal{I}_{2}$, $\mathcal{I}_{-3}$ and $\mathcal{I}_{3}$, $\cdots$, $\mathcal{I}_{-6}$ and $\mathcal{I}_{6}$, until we are sufficiently close to the non-reconnecting interface at $\chi=0$. The distance between each pair of interfaces is equidistant in $\sqrt{\chi}$, so the width of each subregion is similar. Once we have determined the location of the interfaces in terms of VMEC flux label $\psi_{t}$, we can compute the constraints for each SPEC subregion from VMEC output. The constraints $\Delta\psi_{t}$ and $\Delta\psi_{p}$ are computed by taking the difference between their corresponding VMEC values on the neighbouring interfaces. Finally, we need to compute the helicity in each volume. Using (7) and (8), the helicity in VMEC between two flux surfaces $\psi_{t,1}$ and $\psi_{t,2}$ can be computed as
[…]
where $\psi_{p}^{\prime}(\psi_{t})=1/q$ has been used. Consider the gauge convention for SPEC Hudson et al. (2012), the relationship between VMEC helicity and SPEC helicity is given by
$K_{\text{SPEC}}=\left\{\begin{array}[]{cc}K_{\text{VMEC}}+\psi_{t,>}\psi_{p,>}&\text{toroidal region}\\ K_{\text{VMEC}}-\Delta\psi_{p}\psi_{t,<}+\Delta\psi_{t}\psi_{p,<}&\text{annulus region}\end{array}\right.,$ (10)
where $\psi_{t,<}$ and $\psi_{p,<}$ are the corresponding fluxes of the inner interface, while $\psi_{t,>}$ and $\psi_{p,>}$ are those During the “reconnection”, we will remove the interfaces in pairs to reconnect flux surfaces with the same helical flux. After an interface is removed, the constraints for the new combined subregion are recomputed from VMEC given the new interfaces.
@@ III.3 A sequence of equilibria representing the crash process
[…]
@@ IV Comparison with M3D-C1
$\partial_{t}\bm{B}$ =×$=\nabla\times[\bm{v}\times\bm{B}-\eta(\bm{j}-\bm{j}_{\text{eq}})],$ (14)
@@
for the mass density $\rho$, velocity $\bm{v}$, pressure $p$, and magnetic field $\bm{B}$, with the current density $\bm{j}=\nabla\times\bm{B}$ and the adiabatic index $\Gamma=5/3$. The viscous stress tensor is $\bm{\Pi}=-\nu(\nabla\bm{v}+\nabla\bm{v}^{\text{T}})-2(\nu_{\text{c}}-\nu)(\nabla\cdot\bm{v})\bm{I}$ and the heat flux $\bm{q}=-\kappa_{\perp}\nabla(T-T_{\text{eq}})-\kappa_{\parallel}\bm{b}\bm{b}\cdot\nabla T$ with $\bm{b}=\bm{B}/B$ and the temperature $T=Mp/\rho$, where $M$ is the ion mass. The transport coefficients used include uniform mass diffusivity $D=10^{-6}$, isotropic and compressible viscosities $\nu=\nu_{\text{c}}=10^{-6}$ (in the unity of inverse Reynolds number), resistivity $\eta=10^{-6}$ (in the unit of inverse magnetic Reynolds number) and strongly anisotropic parallel and perpendicular thermal conductivities $\kappa_{\parallel}=1$ and $\kappa_{\perp}=10^{-6}$. For this value of $\eta$, the relationship between $\gamma$ and $\eta$ follows $\gamma\sim\eta^{1/3}$ as shown in FIG. 3, and therefore the linear instability is dominated by the resistive kink mode. The equilibrium fields $\rho_{\text{eq}}$, $T_{\text{eq}}$, and $j_{\text{eq}}$ subtracted in the dissipative terms act as effective sources to sustain the equilibrium in the absence of instabilities, and we consider uniform equilibrium density ($\rho_{\text{eq}}=1$) and pressure ($p_{\text{eq}}=10^{-3}$) profiles for simplicity. We use 3807 reduced quintic elements in the $(R,Z)$ plane and 8 Hermite cubic elements in the toroidal direction, and the time step size is $2$. The boundary conditions are ideal on the magnetic field, no-slip on the velocity, and Dirichlet on the density and pressure.
@@
We have selected four different time slices (a) $t=1450$, (b) $t=1550$, (c) $t=1650$, and (d) $t=1750$ (in the unit of $\tau_{A}$). These four slices are close to the four different SPEC equilibria in FIG. 6, namely the 15-, 11- 7-, and 3-subregion equilibria, respectively. The corresponding Poincaré maps are given in FIG. 8. The toroidal plane shown is chosen to match the island phasing in SPEC. They are very close to the SPEC equilibria presented in FIG. 6 by visual comparison. The duration of the crash is $300\tau_{A}$, justifying the assumption that the crash is much slower than the Alfvén time scale $\tau_{A}$ but faster than the resistive time scale $\tau_{R}\sim\eta^{-1}=10^{6}$. The early phase of the crash in FIG. 6 and FIG. 8 (a) and (b) are almost identical. This is because in the early stage, the growth of the island is still slow and the reconnection layer is thin, such that an equilibrium model is more appropriate. There are more notable differences in the later phase of the crash in figures (c) and (d). In SPEC equilibrium (c), the chaotic region only appears around the good flux surfaces surrounding the original and the new magnetic axis, and is bounded by the non-reconnecting interface. This is due to our assumption that the crash only affects plasma within this interface. However, in M3D-C1 the crash is more dramatic and the chaotic region extends beyond the non-reconnecting flux surface to the edge. Finally, SPEC predicts an axisymmetric after-crash configuration, while in M3D-C1, the outer region remains chaotic for a very long time.
@@ V Conclusion
In this study, we have constructed a sequence of Multi-region Relaxed MHD (MRxMHD) equilibria to represent the sawtooth crash process in tokamak plasmas with zero plasma $\beta$. The plasma is partitioned into subregions separated by ideal interfaces, allowing the relaxation locally within reconnected regions while preserving the magnetic topology outside. A helical equilibrium with a magnetic island was identified as having lower energy than the corresponding axisymmetric state. The reconnection region was then progressively enlarged by removing interfaces in pairs, representing the intermediate stages of the sawtooth crash, with the island growing larger and the magnetic energy further reducing. These helical equilibria are all minimum energy states given the constrained topology of the unreconnected region, with the two interfaces bounding the island barely touch each other: the only way to reduce the energy further is via reconnecting these two interfaces. The results demonstrate the capability of MRxMHD in capturing the transition from the initial axisymmetric equilibrium to states with magnetic islands, chaotic regions, and, eventually, the post-crash equilibrium.
Comparison with nonlinear MHD simulations using M3D-C1 reveals qualitative and quantitative agreement: this confirms the sawtooth with low $\beta$ considered in our current paper is indeed a reconnection process. The jump of the $q$ profile matches across the separatrix, showing that MRxMHD is correctly capturing the size of the current sheet. The MRxMHD approach fails to spatially-resolve the poloidal structure of the current sheet due to its coarse assumption of the reconnected region being a single Taylor-relaxed subregion, despite the flat current profile being a good assumption for the majority of the island.

### v2

@@ I Introduction
The main purpose of the current paper is to examine the suitability of MRxMHD in modelling a sawtooth crash as a sequence of equilibria. Sawtooth is a well-known relaxation process in tokamaks von Goeler, Stodiek, and Sauthoff (1974) and current-carying stellarators Nagayama et al. (2003); Zanini et al. (2020). It is triggered when a $q=1$ surface presents inside the plasma volume with the temperature profile undergoing periodic crashes and flattenings, leading to worse confinement in general. Here $q$ denotes the safety factor. The crash is usually followed by a ramp-up recovering phase during which the plasma internal energy re-accumulates. The physical mechanism of the sawtooth is still under debate Chapman (2011). The two most cited models are the Kadomstev model Kadomtsev (1975) and the Wesson model Wesson (1986, 1990). The Kadomstev model attributes the crash to a $m=n=1$ internal kink mode and magnetic reconnection, where $m$ and $n$ are the poloidal and toroidal mode numbers, respectively. During a crash, flux surfaces and their neighbourhood with the same helical flux label $\chi=q_{\text{res}}\psi_{p}-\psi_{t}$ reconnects, conserving $\chi$ itself and the differential toroidal flux. Here $\psi_{p}$ and $\psi_{t}$ are the poloidal flux and toroidal flux, respectively, while $q_{\text{res}}=1$ is the resonant safety factor. The Wesson model, on the other hand, relies on interchange instabilities, which are more prominent at a relatively high plasma $\beta$, where $\beta=2\mu_{0}p/B^{2}$ is the ratio between the plasma kinetic energy and magnetic energy, with $\mu_{0}$ being the vacuum magnetic permeability, $p$ the plasma pressure, and $B$ the magnetic field strength.
@@ II The MRxMHD model
$W_{i}=\int_{\mathcal{R}_{i}}\left(\frac{p}{\Gamma-1}+\frac{B^{2}}{2}\right)dV,$ (1)
@@
$\nabla\times\bm{B}=\mu_{i}\bm{B},\quad\text{within }\mathcal{R}_{i},$ (4)
@@
$\bm{B}\cdot\hat{\bm{n}}=0,\quad\text{on }\mathcal{I}_{i},$ (5)
where $^{\hat{\bm{n}}}$ is the normal unit vector on the interfaces. In the view of the variational principle presented here, $\mu_{i}$ is a dependent variable determined by matching the helicity constraint. Extremising (3) with respect to the location of $\mathcal{I}_{i}$ leads to the force balance condition on each interface given by
@@
where $\left[\left[\cdots\right]\right]$ stands for the difference across the interface. The pressure is then a constant in each subregion, determined by the constraint $p_{i}V_{i}^{\Gamma}=S_{i}$. We emphasise that one needs to specify four constraints for each subregion: $\Delta\psi_{t0,i}$, $\Delta\psi_{p0,i}$, $K_{0,i}$, and $p_{i}$ (or $S_{i}$). Again, for the innermost subregion, the constraint $\Delta\psi_{p0,i}$ is dropped.
@@ III.1 Equilibrium setup
A linear resistive MHD stability calculation with PHONIEX Blokland et al. (2007) shows the equilibrium is unstable to tearing/resistive-kink mode depending on the resistivity. This is demonstrated in FIG. 3 where the linear growth rate $\gamma$ is plotted as a function of plasma resistivity $\eta$. The growth rate $\gamma$ is normalised to the Alfvén frequency on axis $\omega_{A0}=V_{A0}/R_{0}=B_{0}/(R_{0}\sqrt{\mu_{0}\rho_{0}})$, where $\rho_{0}$ is the density on axis and $\mu_{0}$ the plasma vacuum permeability, while the resistivity $\eta$ is in the unit of $\mu_{0}\omega_{A0}R_{0}^{2}$, i.e. the inverse Lundquist number. At low resistivity, the growth rate follows the scaling $\gamma\sim\eta^{3/5}$, a typical relationship for a tearing mode. At a higher resistivity, it gradually transits into a resistive kink mode Coppi et al. (1976) with scaling $\gamma\sim\eta^{1/3}$. Note that the equilibrium is stable to the ideal internal kink mode Bussac et al. (1975) as no instability is found in the limit $\eta\rightarrow 0$ since $\beta=0$, and the cross-section circular.
@@ III.2 Initial MRxMHD equilibrium
Within the reconnection region, we also need to set up a few interfaces. Again, according to the Kadomtsev model, flux surfaces with the same helical flux will reconnect. We therefore place pairs of interfaces at equal helical fluxes based on the VMEC equilibrium. This is demonstrated by FIG. 4. Starting from the subregion containing the $q=1$ surface, we put $\mathcal{I}_{-1}$ inside the $q=1$ surface and $\mathcal{I}_{1}$ outside, such that $\chi(\mathcal{I}_{-1})=\chi(\mathcal{I}_{1})$. Similarly, we place $\mathcal{I}_{-2}$ and $\mathcal{I}_{2}$, $\mathcal{I}_{-3}$ and $\mathcal{I}_{3}$, $\cdots$, $\mathcal{I}_{-6}$ and $\mathcal{I}_{6}$, until we are sufficiently close to the non-reconnecting interface at $\chi=0$. The distance between each pair of interfaces is equidistant in $\sqrt{\chi}$, so the width of each subregion is similar. We note that the final result is not very sensitive to the number of interfaces, as long as they are densely packed in the yet-to-reconnect region and the pairs are chosen to have equal helical flux. Once we have determined the location of the interfaces in terms of VMEC flux label $\psi_{t}$, we can compute the constraints for each SPEC subregion from VMEC output. The constraints $\Delta\psi_{t}$ and $\Delta\psi_{p}$ are computed by taking the difference between their corresponding VMEC values on the neighbouring interfaces. Finally, we need to compute the helicity in each volume. Using (7) and (8), the helicity in VMEC between two flux surfaces $\psi_{t,1}$ and $\psi_{t,2}$ can be computed as
$K_{\text{VMEC}}=2\pi^{2}\int_{\psi_{t,1}}^{\psi_{t,2}}\left(\frac{\psi_{t}}{q}-\psi_{p}\right)d\psi_{t},$ (9)
where $\psi_{p}^{\prime}(\psi_{t})=1/q$ has been used. Considering the gauge convention for SPEC Hudson et al. (2012), the relationship between VMEC helicity and SPEC helicity is given by
$K_{\text{SPEC}}=\left\{\begin{array}[]{cc}K_{\text{VMEC}}+2\pi^{2}\psi_{t,>}\psi_{p,>}&\text{toroidal region}\\ K_{\text{VMEC}}-2\pi^{2}(\Delta\psi_{p}\psi_{t,<}+\Delta\psi_{t}\psi_{p,<})&\text{annulus region}\end{array}\right$., (10)
where $\psi_{t,<}$ and $\psi_{p,<}$ are the corresponding fluxes at the inner interface, while $\psi_{t,>}$ and $\psi_{p,>}$ are those at the outer interface. We label the enclosed flux by $\Delta\psi_{t}=\psi_{t,>}-\psi_{t,<}$ and $\Delta\psi_{p}=\psi_{p,>}-\psi_{p,<}$. During the “reconnection”, we will remove the interfaces in pairs to reconnect flux surfaces with the same helical flux. After an interface is removed, the constraints for the new combined subregion are recomputed from VMEC given the new interfaces.
@@ III.3 A sequence of equilibria representing the crash process
In FIG. 7, we plot the total change of magnetic energy $\int dVB^{2}/2$ of the entire plasma as a function of reconnected helical flux $\Delta\chi$, for both the axisymmetric solution and the non-axisymmetric equilibrium as a percentage of the axisymmetric 15-subregion equilibrium. The non-axisymmetric equilibrium has a lower MHD energy than the corresponding axisymmetric equilibrium. The difference in energy is very small (on the order of 0.01%), even though the magnetic topology changes dramatically, consistent with the finding of Cooper et al. Cooper et al. (2010). It is noteworthy that even when the equilibrium remains axisymmetric as we combine subregions, the total energy will drop. This is because we are reducing the number of constraints (interfaces) which will naturally lead to a lower total energy.
@@ IV Comparison with M3D-C1
$\partial_{t}\bm{B}$ $=\nabla\times[\bm{v}\times\bm{B}-\eta(\bm{j}-\bm{j}_{\text{eq}})],$ (14)
@@
for the mass density $\rho$, velocity $\bm{v}$, pressure $p$, and magnetic field $\bm{B}$, with the current density $\bm{j}=\nabla\times\bm{B}$ and the adiabatic index $\Gamma=5/3$. The viscous stress tensor is $\bm{\Pi}=-\nu(\nabla\bm{v}+\nabla\bm{v}^{\text{T}})-2(\nu_{\text{c}}-\nu)(\nabla\cdot\bm{v})\bm{I}$ and the heat flux $\bm{q}=-\kappa_{\perp}\nabla(T-T_{\text{eq}})-\kappa_{\parallel}\bm{b}\bm{b}\cdot\nabla T$ with $\bm{b}=\bm{B}/B$ and the temperature $T=Mp/\rho$, where $M$ is the ion mass. The transport coefficients used include uniform mass diffusivity $D=10^{-6}$, isotropic and compressible viscosities $\nu=\nu_{\text{c}}=10^{-6}$ (in the unity of inverse Reynolds number), resistivity $\eta=10^{-6}$ (in the unit of inverse Lundquist number) and strongly anisotropic parallel and perpendicular thermal conductivities $\kappa_{\parallel}=1$ and $\kappa_{\perp}=10^{-6}$. For this value of $\eta$, the relationship between $\gamma$ and $\eta$ follows $\gamma\sim\eta^{1/3}$ as shown in FIG. 3, and therefore the linear instability is dominated by the resistive kink mode. The equilibrium fields $\rho_{\text{eq}}$, $T_{\text{eq}}$, and $j_{\text{eq}}$ subtracted in the dissipative terms act as effective sources to sustain the equilibrium in the absence of instabilities, and we consider uniform equilibrium density ($\rho_{\text{eq}}=1$) and pressure ($p_{\text{eq}}=10^{-3}$) profiles for simplicity. We use 3807 reduced quintic elements in the $(R,Z)$ plane and 8 Hermite cubic elements in the toroidal direction, and the time step size is $2$. The boundary conditions are ideal on the magnetic field, no-slip on the velocity, and Dirichlet on the density and pressure.
@@
We have selected four different time slices (a) $t=1450$, (b) $t=1550$, (c) $t=1650$, and (d) $t=1750$ (in the unit of $\tau_{A}$). The size of the reconnection region in these four slices is visually close to the four different SPEC equilibria in FIG. 6, namely the 15-, 11- 7-, and 3-subregion equilibria, respectively. The corresponding Poincaré maps are given in FIG. 8. The toroidal plane shown is chosen to match the island phasing in SPEC. They are very close to the SPEC equilibria presented in FIG. 6 by visual comparison. The duration of the crash is $300\tau_{A}$, justifying the assumption that the crash is much slower than the Alfvén time scale $\tau_{A}$ but faster than the resistive time scale $\tau_{R}\sim\eta^{-1}=10^{6}$. The early phase of the crash in FIG. 6 and FIG. 8 (a) and (b) are almost identical. This is because in the early stage, the growth of the island is still slow and the reconnection layer is thin, such that an equilibrium model is more appropriate. There are more notable differences in the later phase of the crash in figures (c) and (d). In SPEC equilibrium (c), the chaotic region only appears around the good flux surfaces surrounding the original and the new magnetic axis, and is bounded by the non-reconnecting interface. This is due to our assumption that the crash only affects plasma within this interface. However, in M3D-C1 the crash is more dramatic and the chaotic region extends beyond the non-reconnecting flux surface to the edge. In future work, we can remove this interface to allow more relaxation, which might give a better match to M3D-C1. Finally, SPEC predicts an axisymmetric after-crash configuration, while in M3D-C1, the outer region remains chaotic for a very long time.
@@ V Conclusion
In this study, we have constructed a sequence of Multi-region Relaxed MHD (MRxMHD) equilibria to represent the sawtooth crash process in tokamak plasmas with zero plasma $\beta$. The plasma is partitioned into subregions separated by ideal interfaces, allowing the relaxation locally within reconnected regions while preserving the magnetic topology outside. A helical equilibrium with a magnetic island was identified as having lower energy than the corresponding axisymmetric state. The reconnection region was then progressively enlarged by removing interfaces in pairs, representing the intermediate stages of the sawtooth crash, with the island growing larger and the magnetic energy further reducing. These helical equilibria are all minimum energy states given the constrained topology of the unreconnected region, with the two interfaces bounding the island barely touch each other: the only way to reduce the energy further is via reconnecting these two interfaces. The results demonstrate the capability of MRxMHD in capturing the transition from the initial axisymmetric equilibrium to states with magnetic islands, chaotic regions, and, eventually, the post-crash equilibrium. The MRxMHD approach does not of course capture the time dependence of the transition.
Comparison with nonlinear MHD simulations using M3D-C1 reveals qualitative and quantitative agreement: this confirms the sawtooth with low $\beta$ considered in our current paper is indeed a reconnection process. We demonstrated that each state along the path of reconnection is a minimum energy state with the reconnected region being (nearly) Taylor-relaxed and the crescent flux surfaces self-emerging. The jump of the $q$ profile matches across the separatrix, showing that MRxMHD is correctly capturing the size of the current sheet. The MRxMHD approach fails to spatially-resolve the poloidal structure of the current sheet due to its coarse assumption of the reconnected region being a single Taylor-relaxed subregion, despite the flat current profile being a good assumption for the majority of the island. Combined, these results demonstrate that MRxMHD approach is able to model the path of reconnection, not just the post-crash state. The construction of SPEC equilibria from a given VMEC equilibrium is fully automated once the number of subregions is specified. Each SPEC equilibrium typically requires approximately 5 hours of wall-clock time, using a number of CPUs equal to the number of subregions. For comparison, the M3D-C1 simulation presented in this work required a total of 5,000 CPU hours. This highlights the computational efficiency of the MRxMHD approach, which enables the direct construction of intermediate states during the crash.